### v1

Abstract
We give an overview of the 2025 Computational Geometry Challenge targeting the problem Minimum Non-Obtuse Triangulation: Given a planar straight-line graph $G$ in the plane, defined by a set of points in the plane (representing vertices) and a set of non-crossing line segments connecting them (representing edges); the objective is to find a feasible non-obtuse triangulation that uses a minimum number of Steiner points. If no triangulation without obtuse triangles is found, the secondary objective is to minimize the number of obtuse triangles in the triangulation.
keywords:
Computational Geometry, geometric optimization, triangulation, Algorithm Engineering, contest
1 Introduction
[…]
Subsequently, the CG:SHOP Challenge became an event within the CG Week program, with top performing solutions reported in the Symposium on Computational Geometry (SoCG) proceedings. The schedule for the Challenge was advanced earlier, to give an opportunity for more participation, particularly among students, e.g., as part of course projects. For CG Weeks 2020, 2021, 2022, 2023, 2024 the Challenge problems were Minimum Convex Partition [17, 43, 34, 18], Coordinated Motion Planning [22, 14, 32, 42, 13, 31, 41], Minimum Partition into Plane Subgraphs [23, 10, 38, 12, 39, 26, 37], Minimum Convex Covering [24, 15, 2], and Maximum Polygon Packing [25, 16, 33, 28, 4], respectively.
The seventh edition of the Challenge in 2025 continued this format, leading to contributions in the SoCG proceedings.
2 The Challenge: Minimum Non-Obtuse Triangulations
A suitable contest problem has a number of desirable properties.
  1. 1.
    The problem is of geometric nature.
  2. 2.
    The problem is of general scientific interest and has received previous attention.
  3. 3.
    Optimization problems tend to be more suitable than feasibility problems; in principle, feasibility problems are also possible, but they need to be suitable for sufficiently fine-grained scoring to produce an interesting contest.
  4. 4.
    Computing optimal solutions is difficult for instances of reasonable size.
  5. 5.
    This difficulty is of a fundamental algorithmic nature, and not only due to issues of encoding or access to sophisticated software or hardware.
  6. 6.
    Verifying feasibility of provided solutions is relatively easy.
In this seventh year, a call for suitable problems was communicated in May 2024. In response, a number of interesting problems were proposed for the 2025 Challenge. These were evaluated with respect to difficulty, distinctiveness from previous years, and existing literature and related work. In the end, the Advisory Board selected the chosen problem. Special thanks go to Mikkel Abrahamsen (University of Copenhagen) who suggested this year’s problem, motivated by a rich history in geometry and optimization, including [30] and a wide range of previous work described further down.
2.1 The Problem
The specific problem that formed the basis of the 2025 CG Challenge was the following.
Problem: Minimum Non-Obtuse Triangulation
Given: A planar straight-line graph $G$, given by a set of $n$ points $P=\{p_{1},\ldots,p_{n}\}$ in the plane and a collection $E$ of non-crossing line segments, each connecting two points $p_{i}$ and $p_{j}$.
Goal: Inserting a suitable set $S$ of Steiner points as desired, find a triangulation of the convex hull of $P$ that uses $P\cup S$ as vertices, contains $E$, such that all created triangles are non-obtuse. Among all such triangulations, find one that minimizes the number of Steiner points.
To make this problem more accessible, especially for students, we modified it slightly to make it easier to find any feasible solution at all, and to have some more tractable but non-trivial solutions available, and, thus, a more continuous difficulty spectrum. First, we changed the problem to accept obtuse triangles, splitting the objective into first minimizing the number of obtuse triangles, and then minimizing the number of Steiner points. This makes any triangulation feasible and serves as a starting point for iterative improvement. Second, we introduced a boundary to the triangulation area, which encapsulate the area of interest into a simple polygon. While this may first sound like a complicating constraint, it actually simplifies the problem, as Steiner points placed on the boundary are less of a risk to create further obtuse triangles that propagate through the whole instance.
2.2 Scoring
The scoring function was designed to reflect both solution quality and feasibility, while ensuring that even partial progress was rewarded. Each instance was scored individually, and team scores were computed as the sum over all instances. The scoring distinguished between solutions with and without obtuse triangles:
  1. 1.
    Solutions without obtuse triangles: Any solution without obtuse triangles received a minimum score of $0.5$ for the respective instance. The best solution found by any participant, i.e., the one using the fewest Steiner points, was awarded the maximum score of $1.0$. Let $k_{B}$ be the number of Steiner points in such a best known solution and $k_{Y}$ be the number used by another team’s solution. That solution received a score of: $\frac{1}{2}+\frac{k_{B}}{2k_{Y}},$ truncated at $1.0$ if $k_{Y}=k_{B}$.
  2. 2.
    Solutions with obtuse triangles: Solutions with obtuse triangles started from a base score of $0.5$. The score was reduced exponentially based on the number of obtuse triangles. Specifically, for a solution with $v$ obtuse triangles, the score was: $\frac{1}{2}\cdot 0.97^{v}.$ The number of Steiner points was not considered for infeasible solutions.
This scheme provided a smooth incentive structure: teams could improve their scores either by reducing the number of obtuse triangles or, once feasibility was reached, by minimizing Steiner points.
2.3 Related Work
Historically, triangulations have been crucial for applications ranging from computer graphics and mesh generation to geographic information systems (GIS) and finite element analysis. In computer graphics, triangulating surfaces enables the rendering of three-dimensional objects with high precision and efficiency. In mesh generation, triangulations help in discretizing a continuous geometric space into simpler elements that can be easily processed for simulations and analyses, particularly in engineering and physics.
The importance of non-obtuse triangulations in computational geometry is particularly evident in solving partial differential equations using the finite element method. In this method, the domain is divided into triangles, with non-obtuse triangulations ensuring that the resulting matrices are Stieltjes—symmetric positive definite with non-positive off-diagonal entries. This specific matrix structure enhances the performance of iterative methods, such as block Gauss-Seidel, improving their convergence rates. The geometric constraint of avoiding obtuse angles in triangulations not only stabilizes numerical calculations but also optimizes the efficiency of computational processes, making non-obtuse triangulations a crucial consideration in scientific computing and numerical analysis.
Minimum Non-Obtuse Triangulations have been studied for quite a while, with references in DCG and SoCG going back to the 1980s [5], and connections to finite element analysis. Subsequent work considered bounds on the necessary numbers of Steiner points [6, 7, 8]. More recent work [9] considers an algorithm for Planar Straight Line Graphs (PSLGs). The computational complexity for simple polygons and polygons with holes remains an open question in the field.
2.4 Instances
The generation of appropriate instances is critical for any challenge. Instances that are too easily solved undermine the complexity of the problem, rendering it trivial. Conversely, instances that require significant computational resources for preprocessing or for finding viable solutions may offer an unfair advantage to teams with superior computing facilities. This issue is amplified when the instance set is so large that it exceeds the management capacity of a limited number of computers.
We developed five distinct instance generators to create a candidate set of instances:
ortho
The ortho generator produces instances defined by orthogonal polygons specified solely by their boundaries, without additional points or segments. It begins with a random rectangle and randomly introduces orthogonal indentations or extensions along the boundary until the desired complexity is achieved. See Figure 1(a) for an example.
point-set
The point-set generator creates instances comprising a set of points in the plane, with the convex hull of these points forming the boundary. The points are sampled from structured sets, such as those provided by the TSPLib instances [36]. See Figure 1(b) for an example.
simple-polygon
The simple-polygon generator produces instances defined by a simple polygon that is given as the boundary without any additional points or segments. This generator extends the previous approach by starting with a random tour and applying two-opt swaps to form a simple polygon. See Figure 1(c) for an example.
simple-polygon-exterior
The simple-polygon-exterior generator extends the simple-polygon approach by incorporating additional constraints: points and segments define the polygon, while the exterior, bounded by the convex hull of these points, must also be triangulated. Essentially, it augments the previous generator by adding the convex hull. See Figure 1(d) for an example.
simple-polygon-exterior-20
The simple-polygon-exterior-20 generator is based on the simple-polygon-exterior generator but randomly deletes 20% of the segments, such that the polygon no longer partitions the area into simple polygons. See Figure 1(e) for an example.
The orthogonal instances were designed to be the easiest to solve, as relatively simple strategies exist for placing the Steiner points. The point-set instances offer a straightforward approach to iteratively improving the triangulation given their minimal constraints, although effective strategies for placing Steiner points are less apparent. The most challenging instances were the simple-polygon-exterior-20 cases, which feature unstructured constraints that complicate the application of a unified strategy. For those, we were not even certain that a solution without obtuse triangles exists.
Given that even solving smaller instances by hand proved extremely difficult, we restricted the instances to a maximum of $250$ points and generated many small instances, confident that this level of difficulty would be sufficient. We were also aware that the problem demands a significant amount of exact arithmetic, making it impractical to scale up to larger instances – especially for student teams. Overall, $150$ instances were selected from a larger set of candidates to ensure diverse instances based on various instance features.
(a) ortho_60_c423f527
(b) point-set_40_8cbf31aa
(c) simple-polygon_40_12969fc3
(d) simple-polygon-exterior_40_ca26cbd4
(e) simple-polygon-exterior-20_60_e6f13145
Figure 1: Examples of the five instance types used in the challenge. The boundaries are shown in blue, the constrained segments in red, and the points in black.
2.5 Categories
The contest was run in an Open Class, in which participants could use any computing device, any amount of computing time (within the duration of the contest) and any team composition. In the Junior Class, a team was required to consist largely of participants who were eligible according to the rules of CG:YRF (the Young Researchers Forum of CG Week), defined as not having defended a formal doctorate before 2023. Additionally, a team could include one senior academic advisor, who was allowed to provide guidance and assist with writing the final paper (if the team was selected for inclusion in the proceedings), but all programming tasks had to be performed by the junior team members.
2.6 Server and Timeline
The competition was facilitated through a dedicated server at TU Braunschweig, accessible at https://cgshop.ibr.cs.tu-bs.de/competition/cg-shop-2025/. An initial batch of example instances was made available on July 31, 2024, followed by the release of the final benchmark set on September 30th, 2024. The competition concluded on January 22nd, 2025 (AoE).
Participants were provided with a verification tool as an open-source Python package, available at https://github.com/CG-SHOP/pyutils25. This tool provided detailed information on errors and allowed participants to easily investigate and correct their solutions if the server rejected them. To ensure accurate verification and address potential floating point arithmetic issues, the tool leveraged the CGAL library [40]. It also provided easily usable Python-bindings to some exact arithmetic operations that we considered potentially useful for the participants such that also users without a C++ background could use them.
3 Outcomes
Table 1 shows the results for teams that outperformed the trivial Delaunay Triangulation baseline, ranked by performance. The first team achieved a near-perfect score of $149.813$ out of $150$ and was awarded the first place, while the second team achieved a score of $139.769$ and was recognized as the best junior team. These were the only teams to compute solutions without any obtuse triangles across all instances and were invited to contribute to the 2025 SoCG proceedings:
  1. 1.
    Team Naughty NOTers: Mikkel Abrahamsen, Florestan Brunck, Jacobus Conradi, Benedikt Kolb and André Nusser [1].
  2. 2.
    Team Gwamegi, also recognized as the best junior team: Taehoon Ahn, Jaegun Lee, Byeonguk Kang and Hwi Kim [3].
Both winning teams employed local search strategies that iteratively insert, move, and delete Steiner points, which are then completed into a full triangulation via a constrained Delaunay triangulation. The overall winning team refined these actions with more advanced techniques, including a crossover strategy to merge solutions.
Team scores are detailed in Table 1 and further broken down by instance type in Table 2 and by instance size in Table 3.
Rank Team Name Score # feasible Junior Comment
1 Naughty NOTers 149.813 150
2 Gwamegi 139.769 150
3 KITriangle 85.482 147
4 Obtuse Terminators 80.756 114
5 <anonymous> 61.541 81
6 die-obtuse 55.596 54
7 cheetos 25.973 0
8 Delaunay Baseline 25.935 0 Delaunay Triang. as reference
Table 1: Team rankings with scores and junior team status. Only teams outperforming the baseline are shown. # feasible denotes the number of instances solved without obtuse triangles.
Instance Type ortho point-set simple-poly spoly-ext spoly-ext-20
Naughty NOTers 13.000 39.955 21.955 32.904 42.000
Gwamegi 12.704 37.230 20.984 30.437 38.416
KITriangle 7.888 22.986 12.200 18.815 23.592
Obtuse Terminators 8.701 23.589 13.290 15.498 19.678
<anonymous> 8.628 21.855 9.613 8.679 12.765
die-obtuse 9.473 14.046 11.848 8.708 11.521
cheetos 2.858 7.023 4.415 5.309 6.368
Delaunay Baseline 2.909 6.946 4.404 5.309 6.368
Table 2: Team scores across different instance types. For readability, instance type names have been abbreviated: ortho=ortho, point-set=point-set, simple-poly=simple-polygon, spoly-ext=simple-polygon-exterior, and spoly-ext-20=simple-polygon-exterior-20.
Instance Size 10 20 40 60 80 100 150 250
Naughty NOTers 24.883 24.975 19.000 22.000 13.955 17.000 14.000 14.000
Gwamegi 24.509 23.989 17.874 20.341 13.096 15.553 12.376 12.031
KITriangle 18.100 14.652 10.342 11.826 7.519 9.171 6.766 7.107
Obtuse Terminators 18.285 14.715 10.494 11.039 7.012 7.828 5.862 5.521
<anonymous> 16.095 13.102 8.263 8.102 6.478 3.943 3.190 2.367
die-obtuse 18.802 12.807 7.581 5.552 3.501 3.190 2.052 2.112
cheetos 10.069 7.356 3.318 2.326 1.161 1.212 0.357 0.173
Delaunay Baseline 10.057 7.336 3.318 2.322 1.159 1.212 0.357 0.173
Table 3: Team scores broken down by instance size.
Figure 2 illustrates the performance of the top teams across various instance sizes. The increasing gap between the top two teams with larger instances suggests that the winning team employed a more effective scaling strategy. Similarly, Figure 3 shows performance across instance types. The ortho instances are confirmed to be the easiest, while the simple-polygon-exterior and simple-polygon-exterior-20 instances pose the greatest challenges, with the former appearing slightly more difficult. Although the fifth team (die-obtuse) ranked third on ortho instances, it could not maintain competitive performance on other types. The visible variation in rankings between all instance types except simple-polygon-exterior and simple-polygon-exterior-20 indicates that the instance types actually demand distinct strategies.
Figure 2: Scores of the top teams across different instance sizes. The x-axis represents the number of items in the instance, and the y-axis shows the corresponding score.
Figure 3: Scores of the top teams across different instance types. The x-axis represents the instance type, and the y-axis shows the corresponding score.
4 Conclusions
The 2025 CG:SHOP Challenge motivated a considerable number of teams to engage in extensive optimization studies. The outcomes promise further insight into the underlying, important optimization problem. This demonstrates the importance of considering geometric optimization problems from a practical perspective.

### v2

@@
 This makes any triangulation feasible and serves as a starting point for iterative improvement.
-Second, we introduced a boundary to the triangulation area, which encapsulate the area of interest into a simple polygon.
+Second, we introduced a boundary to the triangulation area, which encapsulates the area of interest into a simple polygon.
 While this may first sound like a complicating constraint, it actually simplifies the problem, as Steiner points placed on the boundary are less of a risk to create further obtuse triangles that propagate through the whole instance.
@@
 Conversely, instances that require significant computational resources for preprocessing or for finding viable solutions may offer an unfair advantage to teams with superior computing facilities.
-This issue is amplified when the instance set is so large that it exceeds the management capacity of a limited number of computers.
+This issue is amplified when the instance set is so large that it exceeds the management capacity when using a limited number of computers.
 We developed five distinct instance generators to create a candidate set of instances:
@@
   1. 1.
-    Team Naughty NOTers: Mikkel Abrahamsen, Florestan Brunck, Jacobus Conradi, Benedikt Kolb and André Nusser [1].
+    Team Naughty NOTers: Mikkel Abrahamsen, Florestan Brunck, Jacobus Conradi, Benedikt Kolbe and André Nusser [1].
   2. 2.